OpenStudy (anonymous):

how to get a order/degree/linearity/unknown function and independent variables of (y')^2-3 y y'+xy=0

OpenStudy (ikram002p):

its partial ^_^

OpenStudy (ikram002p):

since u have xy together

OpenStudy (anonymous):

\(\bf{Order}\) is given by the highest order derivative. In this case, the highest derivative is the first, so the DE is first order. \(\bf{Degree}\) is the highest power of the highest order derivative. \(y'\) is squared, so the degree is 2. \(\bf{Linearity}\) is determined much like the linearity of an algebraic equation. \[f(x)y'-g(x)y=h(x)\] is a linear equation (with order and degree 1 - this is a necessary condition for linearity). Compare with the algebraic equation \[ax+b=c\] The algebraic equation will be non-linear if it contains a function of \(x\). Examples: \[ax^2+b=c\\ a+\frac{b}{x}=c\\ a+bx=ce^x\] Similarly, a differential equation will be non-linear if t contains a function of the derivative or the dependent variable. Examples: \[yy'=0\\ (y')^2=y\\ y'-\sin y=x\] Your equation is non-linear. To determine \(\bf{variable~in/dependence}\): a differential equation can be considered a multivariable function: \[f(x,y,y',...,y^{(n)})=0\] Generally, the first variable is the independent variable, the second is the dependent variable, and every variable listed after the dependent is a derivative of the dependent variable. More info here: http://mathworld.wolfram.com/OrdinaryDifferentialEquation.html